My three daughters were each born 2 years apart. Their combined ages is 63 years. What is the age of the eldest ?
Answer by lenny460(1073)

My three daughters were each born 2 years apart. Their combined ages is 63 years. What is the age of the eldest ?
Let their ages be n, n + 2 and n + 4.
Therefore:
n + n + 2 + n + 4 = 63 years
3n + 6 = 63
3n = 57
n = 19
The eldest age is n + 4
19 + 4 = 23

Answer:
23 years